### v1

# envname-P envname#1 BOAD: Discovering Hierarchical Software Engineering Agents via Bandit Optimization

Abstract
Large language models (LLMs) have shown strong reasoning and coding capabilities, yet they struggle to generalize to real-world software engineering (SWE) problems that are long-horizon and out of distribution. Existing systems often rely on a single agent to handle the entire workflow—interpreting issues, navigating large codebases, and implementing fixes—within one reasoning chain. Such monolithic designs force the model to retain irrelevant context, leading to spurious correlations and poor generalization. Motivated by how human engineers decompose complex problems, we propose structuring SWE agents as orchestrators coordinating specialized sub-agents for sub-tasks such as localization, editing, and validation. The challenge lies in discovering effective hierarchies automatically: as the number of sub-agents grows, the search space becomes combinatorial, and it is difficult to attribute credit to individual sub-agents within a team. We address these challenges by formulating hierarchy discovery as a multi-armed bandit (MAB) problem, where each arm represents a candidate sub-agent and the reward measures its helpfulness when collaborating with others. This framework, termed Bandit Optimization for Agent Design (BOAD), enables efficient exploration of sub-agent designs under limited evaluation budgets. On SWE-bench-Verified, BOAD outperforms single-agent and manually designed multi-agent systems. On SWE-bench-Live, featuring more recent and out-of-distribution issues, our 36B system ranks second on the leaderboard at the time of evaluation, surpassing larger models such as GPT-4 and Claude. These results demonstrate that automatically discovered hierarchical multi-agent systems significantly improve generalization on challenging long-horizon SWE tasks. Code is available at https://github.com/iamxjy/BOAD-SWE-Agent.
1 Introduction
Large language models (LLMs) have achieved remarkable progress in natural language processing [38] and reasoning [21], and are increasingly adopted in solving complex coding problems [58]. Yet solving real-world software engineering (SWE) problems remains challenging [25] for LLMs, particularly for issues that fall outside the training distribution [54]. Despite strong results on SWE-bench-Verified [25], state-of-the-art systems struggle on more recent and out-of-distribution issues in SWE-bench-Live [54].
We hypothesize that one reason current SWE agents generalize poorly is their reliance on a single agent to solve a problem within a single reasoning chain. Solving a software engineering task requires handling multiple sub-tasks—such as locating relevant files, editing code, and running tests—each of which depends only on a subset of information. For example, during code editing, the model only needs to know which line to modify, not how that file was previously located. Retaining such irrelevant context introduces spurious correlations [57; 36], leading models trained on long reasoning chains to overfit to training distributions and degrade on out-of-distribution problems.
Motivated by the need to limit irrelevant context during long reasoning processes, we introduce explicit hierarchical structure to LLM-based agents. This design choice mirrors how human engineers solve complex problems—cognitive science studies show that humans reduce cognitive load by decomposing complex tasks into manageable sub-tasks and integrating their outcomes to form the final solution [32; 37; 33]. Recent work [55] demonstrates that decomposing a complex task into smaller sub-tasks improves LLM performance. In software engineering, such decomposition commonly maps to stages like bug reproduction, fault localization, code modification, and validation [45; 24; 31].
This same idea appears as temporal abstraction in hierarchical reinforcement learning (HRL) [11]: instead of solving problems step by step, the agent delegates control to reusable sub-agents, each defined as a policy for a specific sub-task [40]. An orchestrator coordinates these sub-agents by selecting which one to activate, and each sub-agent runs until it completes its sub-task [40; 19; 10]. Planning at the level of sub-agents shortens the effective decision horizon and abstracts away low-level details, allowing the orchestrator to focus on high-level structure. This abstraction supports generalization because sub-agents capture task-invariant skills—such as repo navigation, information retrieval, or code modification—that can be recomposed across different tasks and contexts, rather than overfitting to task-specific action sequences.
Although hierarchical multi-agent systems for SWE tasks appear promising, prior attempts have achieved only limited success. Existing approaches [8; 34; 14] typically rely on manually constructed systems, where sub-agents are assigned to predefined sub-tasks (e.g., locating faulty files) and coordinated by a human-designed orchestrator. These designs depend heavily on human intuition to decompose tasks and assign agent roles, but such decompositions often fail to align with how LLM-based agents actually reason and behave. Because LLM behavior is difficult to predict, workflows that appear logical to humans can perform poorly in practice (Section 5).
This motivates automating multi-agent design, which removes the need for manual specification and enables the discovery of effective agent designs that may not be obvious to human designers. While prior work has explored optimizing the design of single-agent SWE systems [51], these methods do not address the additional challenges of multi-agent coordination, where agents must interact and share information dynamically. Other approaches optimize fixed, multi-step workflows involving multiple LLM calls [26; 52; 23], but they still rely on predefined pipelines. In contrast, SWE agents operate in open-ended environments where execution order and interactions are not fixed in advance.
Automating multi-agent design introduces two main challenges. First, as the number of sub-agents grows, the design space expands combinatorially, and evaluating each multi-agent design on SWE tasks is far slower than on simpler reasoning benchmarks like multi-hop QA [23]. Second, the evaluation signal is coarse and noisy—success or failure of the multi-agent system reveals little about each sub-agent’s contribution. A team of sub-agents may succeed even if one sub-agent consistently makes mistakes and others compensate, making it difficult to assign credit accurately.
To address these challenges, we take a bottom-up approach: rather than exhaustively searching over all combinations of sub-agents, we first identify promising sub-agents individually and then compose them into a team. This makes the search space grow linearly instead of exponentially with the number of sub-agents. The key difficulty is credit assignment—a sub-agent’s value cannot be measured in isolation since it must collaborate with others to solve an issue (e.g., a localizer alone cannot edit code). Instead of judging success by whether the entire team solves a problem, we estimate each sub-agent’s helpfulness, measuring how much it contributes to solving the problem when in combination with other sub-agents. We use LLM-as-a-judge [29] to assess this contribution from the agent’s behavior during the problem-solving trajectory, yielding a more informative signal than a binary success or failure outcome.
Once helpfulness is defined, the next question is how to efficiently discover effective sub-agents. Evaluating a sub-agent once is insufficient because its performance depends on its teammates and the task instance, both of which vary stochastically. Yet repeatedly testing all sub-agents would be prohibitively expensive. To balance exploration (trying under-tested sub-agents) and exploitation (focusing on those that have performed well), we formulate multi-agent design as a multi-armed bandit (MAB) problem [17; 1; 20; 50], where each arm corresponds to a sub-agent and the reward is its helpfulness. Well-established exploration algorithms in multi-armed bandits (MAB) [28] provide an effective toolkit for adaptively selecting sub-agents to evaluate and efficiently discovering a set of helpful ones. We refer to this framework as Bandit Optimization for Agent Design (BOAD).
We evaluate our BOAD on SWE-bench-Verified [25], a benchmark for software engineering tasks grounded in real GitHub issues. On SWE-bench-Verified, our method consistently outperforms single-agent systems and manually designed multi-agent systems. On SWE-bench-Live [30], which includes more recently collected issues and presents out-of-distribution challenges, our system achieves 3rd place on the leaderboard using a 36B model—outperforming larger-scale systems based on Claude and GPT-4. To our best knowledge, our work is the first to show generalization improvements using automatically discovered hierarchical multi-agent systems on challenging long-horizon interactive tasks like SWE-bench.
2 Related works
Prior works [8; 34; 14] built multi-agent systems for SWE tasks by manually designing the prompts of each sub-agent and the orchestrator that coordinates them. Each sub-agent handled a specific sub-task (e.g., issue localization). While intuitive, these systems required heavy engineering and yielded only marginal gains, leading to limited follow-up work in this direction. Other efforts [30; 53] leveraged diversity across multiple agents or models to improve accuracy through ensembling, where each agent generated a patch and an aggregator selected the final solution. In contrast, we aim to automatically learn a hierarchical multi-agent system—comprising an orchestrator and multiple sub-agents—without manually defining their roles or coordination.
Automating multi-agent design is closely related to prior work on optimizing single-agent systems. Existing approaches have focused on improving an agent’s prompt [26; 2; 22] or its scaffold [51], including both tools and prompts, by iteratively refining these components based on task performance. These methods typically optimize a single agent for a given task. In contrast, we study the automated design of multi-agent systems, where different agents are specialized for distinct sub-tasks and must coordinate to solve the overall problem.
Closer to multi-agent systems, several prior works study how to optimize the workflow among agents. In these approaches, a workflow may be represented as a Python function that orchestrates multiple LLM calls Hu et al. [23], or as a graph of agents with distinct roles and prompts [52; 56]. While effective in their respective settings, such workflow optimization methods have seen limited adoption in SWE tasks. One likely reason is that SWE tasks often demand fine-grained, dynamic coordination that evolves as the task unfolds. In contrast to workflows that are largely specified prior to execution, multi-agent systems for SWE typically require agents to adapt and coordinate on the fly in response to intermediate outcomes and newly discovered information.
Prior work has explored automatically generating multi-agent systems [16; 15; 49], but these approaches typically assume a fixed execution scheme, such as a predefined agent ordering [16; 15] or parallel aggregation of agent outputs [49]. In contrast, our work focuses on optimizing how agents coordinate with one another. In addition, existing studies primarily consider simple, inexpensive tasks, whereas we target SWE tasks that involve long-horizon, multi-turn interactions and are costly to evaluate, motivating the need for higher sample efficiency.
3 Preliminaries
Software engineering agents
We study the problem of using LLMs to resolve real-world GitHub issues, where each issue consists of a textual description and a corresponding code repository. Since issues are not self-contained, solving them requires identifying and modifying relevant parts of the codebase. In this work, we focus exclusively on agentic methods [48], where an LLM interacts with a runtime environment through tool use. Such agents can browse files, execute shell commands, run tests, and edit code directly, giving them the flexibility to tackle long-horizon tasks end-to-end.
Markov Decision Process (MDP)
We model agent–environment interaction as a finite-horizon Markov decision process (MDP) [35], . At each step , the agent observes a state , consisting of the issue description and the history of prior tool interactions . The agent samples an action from its policy , where includes all available tools and commands. Executing yields an observation (e.g., logs, diffs, or test results), updating the state to . A trajectory terminates at when the agent submits a patch (forced at if none is submitted earlier). Rewards are sparse:
The agent’s goal is to maximize the expected rewards . With sparse rewards and long horizons, discovering successful trajectories is challenging.
Temporal Abstraction via Semi-MDP (SMDP)
Semi-Markov Decision Process (SMDP) framework [39] is widely used to mitigate long-horizon sparse reward problems. Instead of issuing primitive actions , the orchestrator selects a temporally extended action (option) . Each option corresponds to a sub-agent that executes a sequence of actions until termination, after which control returns at , where denotes the duration of an sub-agent. This reduces decision frequency and simplifies planning.
Multi-Armed Bandit (MAB)
Multi-armed bandit (MAB) [28] is a special case of an MDP with a single state and no transitions. At each round , the learner selects an arm , receives a stochastic reward drawn from an unknown distribution, and seeks to maximize the cumulative reward over a fixed interaction budget . The MAB framework captures decision-making under uncertainty when only a limited number of trials are available.
4 Method: Bandit Optimization for Agent Design (BOAD)
Our goal is to automatically discover a set of sub-agents together with an orchestrator that maximizes expected task performance (i.e., reward) . A straightforward approach is to directly optimize over both components using evolutionary [13] or random search [12]:
where denotes a design set of representative problem instances. However, evaluating a candidate pair requires executing full interaction trajectories and repeatedly querying the reward function . In software engineering (SWE) tasks, such evaluations are long-horizon, multi-turn, and computationally expensive, making naive joint search over impractical.
Agent design as a multi-armed bandit. To address this challenge, we reformulate agent discovery as a multi-armed bandit (MAB) problem [28]. In this formulation, each arm corresponds to a candidate sub-agent , and at each round , a subset of sub-agents is selected and evaluated under an orchestrator. Casting agent design as a bandit problem allows us to leverage principled exploration–exploitation strategies [27; 9]: these algorithms can allocate more evaluations to promising sub-agents while still exploring alternatives. This substantially reduces wasted computation on poor designs and makes automatic discovery of multi-agent systems feasible under the high evaluation costs characteristic of SWE tasks. We present the formal bandit formulation in Section 4.1. Towards this end, we must solve the following problems:
1.
The space of possible orchestrators and sub-agent sets is vast and initially unknown, making it infeasible to pre-enumerate arms, as required by standard bandit algorithms.
2.
Even if we evaluate a sub-agent along with an orchestrator and the other sub-agents, credit assignment is ambiguous: some sub-agents may succeed only by “free-riding” on others, so the observed reward does not necessarily reflect their individual contribution.
Next, we illustrate how we tackle these challenges in the following sections.
4.1 Agent Design as a Multi-Armed Bandit Problem
The space of orchestrator–subagent pairs is vast and infeasible to enumerate. To make the search tractable, we maintain an archive of candidate sub-agents. Instead of treating each subset of sub-agents as an arm, we treat each sub-agent as an arm, enabling information sharing across different sub-agents (will be explained below). At each round , the algorithm selects a subset by choosing arms, instantiates an orchestrator for , evaluates on example problems from a design set, and propagates feedback to all participating sub-agents. Feedback is calculated for each sub-agent independently to tackle the credit assignment problem (details in 4.2). Because sub-agents appear in multiple subsets, each evaluation, even if unsuccessful, provides signal for multiple subsets at the same time. This formulation supports efficient credit assignment, reduces redundant exploration, and progressively refines estimates for each . Algorithm 1 summarizes the procedure, with further details provided below.
Bootstrapping a sub-agent archive
We begin with an initial archive of candidate sub-agents. This archive is generated by prompting an LLM with the template in Appendix A.1.2. However, simply generating sub-agents is insufficient because the orchestrator may not know how to invoke them. We present sub-agents as tools to the orchestrator and adopt the standard tool-calling protocol from SWE-agent [48]. To call a sub-agent, the orchestrator must parse its docstring to understand the functionality and supply the required inputs. For example, an issue-localizer sub-agent requires the issue summary as input; without it, the sub-agent cannot operate. To ensure this, we introduce a warm-up stage that rewrites each generated sub-agent’s docstring into a precise specification of its inputs and outputs, enabling the orchestrator to integrate it correctly. Details are provided in Appendix A.1.1.
Sub-agent evaluation
At each round , we select a set of sub-agents . Given this set, an orchestrator is instantiated by prompting an LLM (see Appendix A.1.3), and the system is evaluated on a subset of example problems from a design set. This evaluation yields a performance score for each sub-agent . A straightforward choice for is the success rate of the system on the design set, but as discussed in Section 4.2, this metric is suboptimal and we propose a more effective alternative.
Balancing exploration and exploitation on sub-agent selection
After bootstrapping the archive, the next challenge is deciding which sub-agents to evaluate in each round. To balance exploration with exploitation, we adopt the Upper Confidence Bound (UCB) [28] strategy. For each sub-agent , we track its empirical mean of the performance score of and selection count up to round . The UCB score of a sub-agent at round is defined as
The first term favors sub-agents with high observed performance, while the second term gives an optimism bonus to under-sampled sub-agents (i.e., exploration). At each round , we select the top- sub-agents based on their UCB scores, ensuring that evaluations increasingly focus on strong candidates while still allocating time to uncertain ones.
Expanding the archive
A fixed archive risks stagnation: once UCB identifies a few strong sub-agents, it will repeatedly exploit them, leaving little opportunity to discover new behaviors. To address this, we expand the archive dynamically using a Chinese Restaurant Process (CRP) [3; 44]. At each round , we prompt the LLM to generate a new sub-agent distinct from those in the current archive (see Appendix A.1.2). The probability of introducing a new sub-agent is
where is a concentration parameter. This mechanism ensures diversity: when the archive is small, new sub-agents are frequently added; as the archive grows, the probability decreases, shifting the emphasis toward reuse of existing ones. Over time, the expected number of distinct sub-agents after rounds grows as , providing unbounded but controlled expansion. We also run the warmup stage (Appendix A.1.1) to ensure the sub-agent is usable by the orchestrator.
4.2 Hindsight Credit Assignment
A central challenge in our framework is defining the performance score of individual sub-agents . A simple approach is to set the score of a sub-agent to the success rate of all trajectories that include it:
where is the set of trajectories at round in which is used by the orchestrator . However, this suffers from a “free-rider” problem: a sub-agent may appear effective simply because it often co-occurs with strong sub-agents, even if it contributes little itself.
To overcome this, we adopt a hindsight-based credit assignment strategy. The idea is to reward a sub-agent whenever its actions help the orchestrator make progress toward solving the problem, even if the orchestrator ultimately fails. Thus, sub-agents that provide useful intermediate steps are credited, while those that do not are penalized, regardless of the outcomes. Concretely, let denote the trajectory of actions and observations produced during problem solving. For each sub-agent that appears in , we query an LLM judge (Appendix A.1.4) with the trajectory and obtain a binary label where indicates that the LLM judge deems ’s contribution in the trajectory as helpful. The performance score of sub-agent is then defined as the empirical average over all evaluated trajectories:
This hindsight-based score provides a more reliable estimate of the utility of than success rates. By directly linking credit to judged contributions, it avoids free-riding effects.
5 Experiments
Our experiments address the central question: Can properly designed hierarchical multi-agent systems improve the generalization performance of SWE agents? We further analyze how the systems discovered by our algorithm differ from human-designed ones and examine the contribution of each design choice to the overall performance gains.
5.1 Setup
Task format and datasets.
We evaluate on the SWE-bench benchmarks: SWE-bench Verified (500 instances) [25] and SWE-bench Live (300 instances) [54]. Verified is a curated, frozen set of real GitHub issues, while Live continuously adds newly collected, human-verified issues from active repositories, making it more resistant to data contamination [46] and better suited for testing generalization to out-of-distribution problems. Each instance includes a GitHub issue, a repository-specific container image, and an executable test harness. The agent must interact with the repository (files and, when available, history) and produce a patch that resolves the issue by passing all tests (pass-to-pass and fail-to-pass).
To avoid overfitting and limit design-time compute, we construct a small yet diverse design set by sampling one random issue per repository (12 total) from Verified, which we find to be sufficient based on an ablation over design-set size (A.3.1). The design set is disjoint from all issues in Live. All results in Tables 2 and 3 are reported on Verified and Live (lite) splits. We also report the result of BOAD on Verified (held out) that exclude the 12 issues used in the design set.
Optimization Details
During BOAD optimization, we run rounds of the bandit loop (Algorithm 1) (testing on up to 100 rounds shows that sub-agents created after around 20 rounds are generally worse, and 20 iterations is enough to converge to a state where helpful rate and UCB rankings align). Each time a new sub-agent is created, the sub-agent first goes through a warm-up process, which uses randomly sampled instances from the design set to iteratively refine the documentation/instance prompt of the sub-agent, ensuring that the sub-agent is usable. We use rounds for the warm-up process. Each round samples sub-agents and evaluates them on the design set. The optimization took around 12 hours on a machine with 56 CPU cores and 440GB RAM. For running SWE-agent (orchestrator and sub-agent LLM inference), we deploy Seed-OSS-36B-Instruct on one H100 GPU node. Note that by the tenth iteration, the top-2 most helpful subagents are the same as the ones converged on later. Running iterations took less than hours.
Implementation.
All experiments use the SWE-agent scaffold with a set of default tools from SWE-agent [48]: edit_anthropic (file viewing/editing), bash (restricted shell commands), and submit. The orchestrator calls sub-agents through the same XML-based tool-calling API, passing information via a context parameter; sub-agents return outputs through this channel without access to the orchestrator’s history. We use Claude-4 to generate candidate designs (Section 4.1) with temperature and evaluate sub-agent helpfulness (Section 4.2). For execution, both orchestrator and sub-agents use Seed-OSS-36B-Instruct with temperature , unless specified, a strong instruction-following model that is not heavily tuned on SWE tasks. This choice ensures improvements reflect the benefit of orchestration rather than fine-tuning on SWE-task specific data. Each sub-agent is equipped with prompts discovered by BOAD, defined with docstrings and argument specs, and given access to the same default tools as the orchestrator.
Finally, for evaluation, we use the two subagents with the highest helpfulness score (the hindsight-based score) found during the subagent discovery (Appendix A.2.1). We also ablate by varying k in the top-k selection and by using success rate instead of helpfulness as the ranking metric (Section 5.3).
5.2 Main results
Success Rate
Table 1 shows that with Seed-OSS-36B-Instruct, BOAD resolves 20.0% of issues on Live, ranking second on the leaderboard at the time of evaluation and outperforming larger models in popular scaffolds (e.g., GPT-4o, DeepSeek-V3, GLM-4.5-Air, Claude 3.7 Sonnet). This is a 63% improvement over the same model with default SWE-agent tools. On Verified, BOAD achieves 53.12%, surpassing many larger models (e.g., GPT-4o, Claude 3.5 Sonnet OpenHands, DeepSeek-R1, DeepSeek-V3) and setting a new state of the art among smaller models (e.g., Qwen3-Coder-30B-A3B-Instruct, Devstral-Small-2505), with a 13.4% gain over the default SWE-agent. Interestingly, adding manually designed sub-agents (Appendix A.2.4) from prior work [8; 34; 14] lowers performance, indicating that human-crafted roles can be misaligned with LLM behavior. Overall, these results demonstrate that BOAD automatically discovers orchestrator–sub-agent structures that not only boost in-distribution performance but also generalize more effectively to out-of-distribution tasks.
Comparison with Evolutionary Baseline
We also implement an evolutionary search for a multi-agent system adapted from Automated Design of Agentic Systems [23], as a baseline to compare BOAD against. Implementation details for the evolutionary search are provided in Appendix A.4.1. When using the same number of evaluation instances (i.e number of SWE-Bench patches generated), we find that the sub-agents generated by the evolutionary search achieve worse performance (17.0% vs 20.0% on SWE-Bench-Live) than BOAD. Additionally, for the same number of iterations, the cost of Claude API calls is more than double than that of BOAD due to the need of generating many sub-agents at each iteration, whereas BOAD reuses subagents across iterations.
| Scale | Model | Scaffold | Verified Resolved (%) | Live Resolved (%) |
| Large | GPT-4o [48] | SWE-agent | 23.0 | 10.0 |
| GPT-4o [47] | Agentless | 38.8 | 11.7 | |
| Claude 3.5 Sonnet [5] | Agentless | 50.8 | – | |
| Claude 3.5 Sonnet [5] | OpenHands | 53.0 | – | |
| Claude 3.7 Sonnet [4] | SWE-agent | 62.4 | 13.7111The SWE-bench-live leaderboard score was 17.7, based on an earlier issue set from April 2025. | |
| Claude 4.0 Sonnet [6] | SWE-agent | 66.8 | – | |
| Claude 4.0 Sonnet [6] | OpenHands | 70.4 | – | |
| DeepSeek-R1 [21] | Agentless | 49.2 | – | |
| DeepSeek-V3 [18] | Agentless | 42.0 | 13.3 | |
| GLM-4.5-Air [41] | OpenHands | 57.6 | – | |
| GLM-4.5-Air [41] | SWE-Agent | – | 17.7 | |
| Qwen3-Coder 480B/A35B Instruct [43] | OpenHands | 69.6 | 24.7 | |
| Small | Qwen3-Coder-30B-A3B-Instruct [43] | SWE-agent | – | 17.0 |
| Qwen3-Coder-30B-A3B-Instruct [43] | OpenHands | 51.6 | – | |
| Devstral-Small-2505 [7] | OpenHands | 46.8 | – | |
| Seed-OSS-36B-Instruct [42] | SWE-agent (baseline) | 49.8 | 12.3 | |
| Seed-OSS-36B-Instruct [42] | SWE-agent + Manual Sub-agent | 47.4 | 14.0 | |
| Seed-OSS-36B-Instruct [42] | SWE-agent + Evolutionary Search | 46.0 | 17.0 | |
| Seed-OSS-36B-Instruct [42] | SWE-agent + BOAD | 53.222253.1 on the SWE-bench-verified set excluding the 12 issues used in the design set. | 20.0 |
Token Analysis
In addition to success rate, we analyze the test-time token usage of the hierarchical multi-agent system discovered by BOAD in comparison to the default single-agent system. Hierarchical multi-agent systems introduce communication overhead, since agents must exchange information, but they can also reduce context length: sub-agents focus on specialized sub-tasks while the orchestrator handles high-level coordination without low-level details. Table 2 compares token usage between SWE-agent and BOAD. Total tokens refer to the average sum of input and output tokens per issue, while max input tokens capture the average maximum input length per instance. Surprisingly, the total token count is comparable—and even lower on SWE-bench-live—than in the original SWE-agent. Moreover, BOAD consistently reduces input tokens, confirming that task decomposition shortens context length.
| Metric | Setting | Verified | Live |
| Total tokens (M) | SWE-agent | ||
| SWE-agent + BOAD | (+0.7%) | (-23.8%) | |
| Max input tokens | SWE-agent | k | k |
| SWE-agent+BOAD | k (-11.6%) | k (-25.0%) |
5.3 Ablation Studies and Analysis
| Research Question | Configuration | SWE-Bench Live (%) |
| Does prompt optimization explain the gains? | w/o Sub-agent | 16.3 |
| w Sub-agent | 20.0 | |
| Do more sub-agents improve performance? | Top-5 sub-agents | 13.7 |
| Top-4 sub-agents | 16.7 | |
| Top-3 sub-agents | 16.3 | |
| Top-2 sub-agents | 20.0 | |
| Top-1 sub-agent | 16.3 | |
| Do we need to customize the orchestrator? | w/o customization | 16.7 |
| w customization | 20.0 | |
| Is expanding the sub-agent archive needed? | w/o expansion | 17.0 |
| w expansion | 20.0 | |
| Is hindsight credit assignment necessary? | Top-3 subagents (success rate) | 11.3 |
| Top-3 subagents (helpfulness) | 16.3 | |
| Top-2 subagents (success rate) | 15.3 | |
| Top-2 subagents (helpfulness) | 20.0 | |
| Are discovered sub-agents transferable to other models? | Claude 3.7 Sonnet | 13.7 |
| + Top-2 sub-agents (helpfulness) | 16.3 |
Does prompt optimization explain the gains? One possible explanation for BOAD ’s performance improvement is that it simply arises from better prompt optimization of SWE-agent. To test this, we introduce a baseline that optimizes the SWE-agent prompt without adding sub-agents (w/o Sub-agent). We run 10 iterations: in each, a new SWE-agent prompt is generated by prompting Claude-4 with the template shown in A.1.5, evaluated on 12 issues (the same setting as BOAD). The first iteration is initialized without history, and from the second onward, prompt generation is conditioned on the top five prompts from previous rounds, ranked by performance. Results in Table 3 show that prompt optimization alone does not reach the performance of BOAD, indicating that the gains are not solely due to prompt tuning but from the discovery of effective sub-agents and orchestration.
Do more sub-agents improve performance? One might expect performance to improve as more sub-agents are added, since each can specialize. To test this, we vary the number of top- sub-agents from 1 to 5 based on the helpfulness score (Section 4.2) and evaluate on Live. Surprisingly, Table 3 shows that performance peaks with exactly two sub-agents, achieving 60/300 (20.0%). A single sub-agent (49/300) fails to leverage specialization, while larger teams of three (49/300), four (50/300), or five (41/300) reduce performance due to communication and coordination overhead. These results suggest that small, focused teams strike the best balance, outperforming both minimal and overly large teams of sub-agents.
Do we need to customize the orchestrator? We next ask whether gains come solely from sub-agent discovery or if the orchestrator must also adapt to its team. We compare two prompting strategies: (i) a generic prompt encouraging sub-agent calls (Appendix A.1.1), and (ii) a customized prompt generated by Claude-4 that explicitly references the top two sub-agents (selected by helpfulness scores) and outlines a plan for using them. Both settings use the same sub-agent set, but only the customized prompt allows the orchestrator to reason about and plan calls to specific sub-agents. Results in Table 3 show the customized orchestrator (w customization) achieves 60/300 (20.0%), versus 50/300 (16.7%) (w/o customization) for the generic one. This indicates that while sub-agents provide new capabilities, the orchestrator must also be specialized to effectively coordinate them.
Is expanding the sub-agent archive needed? As discussed in Section 4.1, the initial archive may be limited, and adding new sub-agents during the design process could be necessary to discover stronger ones. To test this, we compare orchestrator performance using (i) sub-agents from the initial archive (w/o expansion) and (ii) sub-agents selected at the end of the design process (w expansion). Both settings use two sub-agents, consistent with our best configuration in Section 5.2, and the orchestrator is generated as described in Section 4.1. Results in Table 3 show that final sub-agents outperform those from the initial archive, highlighting the importance of expanding the archive over time.
Is hindsight credit assignment necessary? To address free-riding issues in sub-agent selection (Section 4.1), we use a helpfulness score to measure each sub-agent’s contribution. To test its importance, we compare orchestrator performance when sub-agents are selected by (i) individual success rate versus (ii) helpfulness score. As shown in Table 3, helpfulness-based selection consistently outperforms success-rate selection, indicating that hindsight credit assignment (Section 4.2) is essential for identifying useful sub-agents.
Are discovered sub-agents transferable to other models? Since BOAD optimizes sub-agents for a specific model, we ask whether the best sub-agents differ across models and whether effective sub-agents can transfer. To test this, we apply the sub-agents from Section 5.2 to SWE-agent+Claude-3.7-Sonnet. As shown in Table 3, the discovered sub-agents do transfer to some extent, though the gains are smaller than those achieved with Seed-OSS-36B-Instruct, the model used for sub-agent optimization.
5.4 Qualitative Analysis of Single- vs Multi-Agent Outcomes
We manually inspected trajectories in which the single- and multi-agent systems produced different outcomes. Three recurring patterns emerged:
1.
Over-editing (multi-agent advantage). Single agents frequently produced extremely long patches, including attempts to create new tests and edits outside the scope of the bug. Such patches inflate apply time and increase the chance of failing pass-to-pass, even if the agent is able to address the primary fault. In contrast, the multi-agent system tended to emit short, localized patches, highlighting the advantage of separating phases like localization and editing/testing.
2.
Multi-site fixes and coverage (multi-agent advantage). When the fix required edits at multiple call sites or modules, the single agent often either over-edited unrelated regions or missed one or more necessary locations. The hierarchical system mitigated both omission and extraneous edits by delegating to a sub-agent for localization, which did a thorough analysis of the repository before making any targetted modifications.
3.
Error propagation from unvalidated sub-agent outputs (multi-agent failure mode). In a minority of cases, the multi-agent system failed while the single agent succeeded. Inspecting these outputs, we found that erroneous sub-agent outputs (e.g., incomplete span identification, misinterpretation of the issue) were accepted as ground truth by the orchestrator, leading subsequent steps astray. Because there is no intermediate validation or self-checking, the orchestrator has limited ability to recover from such upstream.
These observations align with our hypothesis: hierarchical delegation constrains edit scope and improves coverage of multi-site fixes, but introduces a new dependency on the quality of sub-agent handoffs. Incorporating lightweight verification (e.g., span cross-checks, invariant tests, or dual-read localization) is a promising mitigation for the third failure mode.
6 Discussion & Conclusion
We present BOAD, a framework that formulates hierarchical multi-agent design as a sequential, online decision making problem to automatically discover multi-agent systems for long-horizon software engineering tasks. Our experiments show that automatically discovered sub-agents, when combined with a customized orchestrator, outperform single-agent and manually designed multi-agent systems on both SWE-Bench Verified and SWE-Bench Live.
Limitations and Future Work: We find that discovered sub-agents transfer across models only partially and failure cases highlight error propagation when orchestrators unconditionally accept sub-agent outputs. Future work should explore evolution on large models, adaptive team sizing, verification, as well as extending the framework to domains beyond software engineering.
Acknowledgment
We thank members of the Improbable AI Lab at MIT and MIT-IBM Watson AI Lab for helpful discussions and feedback.
Author Contribution
Iris Xu: Led the project, implemented and designed the core algorithm, conducted all the experiments, and wrote the experiment section.
Guangtao Zeng: Implemented the baselines, explored the preliminary design of the ideas of this work, and wrote the related work section.
Zexue He: Explored the preliminary design of the ideas of this work and helped revise the manuscript.
Charles Jin: Assisted and supported the implementation of our method.
Aldo Pareja: Assisted and supported writing.
Dan Gutfreund: Coordinated the compute and infrastructure.
Chuang Gan: Coordinated the compute and infrastructure.
Zhang-Wei Hong: Led the writing and ideation of the project, coordinated the team and resources, and explored the preliminary design of the ideas of this work.
Ethics Statement
Coding agents hold strong promise for automating code generation and bug fixing, but they also carry risks of unintended or harmful outputs. For instance, an LLM-based agent may produce commands that could compromise a system (e.g., downloading unauthorized packages or deleting user files with rm -rf). To mitigate these risks in our study, all experiments were conducted within Docker containers, providing isolated and sandboxed environments that prevent harmful commands from impacting real user devices and substantially reducing the potential for actual harm. Our implementation also builds upon the SWE-Agent framework, which has been previously published and reviewed under established ethical standards. We carefully follow its curated protocols and licensing requirements.
Nevertheless, as with any AI-based coding agent framework, there remains the risk of deliberate misuse, for example, a malicious user prompting the system to generate harmful or hacking code. While our contribution is centered on advancing the technical design of hierarchical coding agents, we emphasize that real-world deployment should be coupled with responsible auditing and oversight, so that potential misuse and unintended consequences can be effectively mitigated.
Reproducibility Statement
For all open-source LLMs (e.g., Seed-OSS-36B-Instruct, Qwen3-Coder-30B-A3B-Instruct), we rely on their official releases. Commercial LLMs and LLM-based tools are accessed through their official APIs and reference implementations. We provide detailed implementation notes of BOAD, including the prompt design for meta-agents, sub-agents, and LLM judges, in Section 5.1. To further support reproducibility and foster future research, we will also release all of our code, used data, and prompts.
Appendix A Experiment Details
A.1 Prompt templates
A.1.1 Prompt template for refining subagent during warmup stage
A.1.2 Meta agent prompts
A.1.3 Custom orchestrator plan prompt
A.1.4 Prompt template for checking if a subagent was helpful in a given instance
A.1.5 Orchestrator-only Prompt
A.2 Subagents
A.2.1 BOAD Top 2 Discovered Subagent Configurations for Seed-OSS-36B-Instruct
A.2.2 All BOAD Discovered Subagents
Here we report all sub-agents discovered over 100 iterations of BOAD. For each sub-agent, we report the iteration number where it was first proposed (Generated Iteration), the number of times it was selected during optimization (), and its final hindsight helpfulness rate.
| Subagent | Generated Iteration | Helpfulness | |
| code_navigator | 1 | 1140 | 0.933 |
| test_runner | 1 | 120 | 0.625 |
| code_fixer | 1 | 36 | 0.361 |
| fix_validator | 2 | 36 | 0.333 |
| issue_reproducer | 3 | 564 | 0.768 |
| dependency_resolver | 5 | 24 | 0.125 |
| test_analyzer | 6 | 24 | 0.083 |
| issue_analyzer | 7 | 1116 | 0.982 |
| precision_editor | 8 | 120 | 0.642 |
| multi_file_coordinator | 10 | 24 | 0.042 |
| code_detective | 11 | 60 | 0.500 |
| config_manager | 14 | 12 | 0.000 |
| git_resolver | 26 | 12 | 0.000 |
| error_debugger | 32 | 12 | 0.000 |
| api_analyzer | 36 | 24 | 0.208 |
| performance_analyzer | 46 | 36 | 0.250 |
| compatibility_checker | 48 | 36 | 0.333 |
| spec_analyzer | 53 | 36 | 0.361 |
| data_flow_analyzer | 60 | 96 | 0.562 |
| refactor_architect | 96 | 24 | 0.042 |
A.2.3 All Evolution Discovered Subagents
Here we report all bundles discovered over 20 iterations of BOAD (one for each iteration). For each sub-agent, we report its iteration number, the generated subagents, and the bundle’s success rate. Note that sub-agents with the same name across different iterations may have different system/instance templates because sub-agents are not reused. On evaluation, we choose the bundle with the highest success rate (latest bundle if tied).
Sub-agents focused on problem analysis or file localization—such as issue_analyzer (0.968 average helpfulness), code_navigator (0.917), and issue_reproducer (0.817)—consistently appear more useful. Our observation is that these agents provide value independently of how later stages unfold: even if code editing or testing fails, identifying the correct files and clarifying the underlying issue is almost always beneficial.
In contrast, sub-agents whose usefulness depends on earlier steps tend to show lower average helpfulness. For example, test_analyzer (0.167) is only useful after the system has already identified the faulty files and produced a candidate patch; if either prerequisite fails, it cannot contribute, which naturally lowers its average helpfulness.
Specialized sub-agents such as dependency_resolver (0.250) and config_manager (0.000) also have low average helpfulness, largely because dependency or configuration issues appear in only a small fraction of tasks.
Finally, once core analysis is completed and the faulty files are correctly located, the orchestrator can often complete the remaining code-editing steps on its own. This explains why code-editing sub-agents like code_fixer do not exhibit high average helpfulness.
| Iteration | Subagents | Success Rate |
| 1 | issue_localizer, issue_reproducer, patch_editor | 0.333 |
| 2 | fix_validator, fix_planner, code_explorer | 0.500 |
| 3 | issue_localizer, test_runner, code_patcher | 0.417 |
| 4 | repo_mapper, code_searcher, fix_validator | 0.500 |
| 5 | code_editor, issue_reproducer, issue_localizer | 0.182 |
| 6 | fix_validator, code_locator, problem_analyzer | 0.583 |
| 7 | test_runner, code_editor, code_analyzer | 0.583 |
| 8 | code_locator, fix_validator, issue_reproducer | 0.417 |
| 9 | code_editor, repo_analyzer, test_runner | 0.250 |
| 10 | code_locator, issue_reproducer, solution_planner | 0.417 |
| 11 | code_patcher, test_validator, repo_explorer | 0.417 |
| 12 | issue_reproducer, code_locator, fix_implementer | 0.364 |
| 13 | fix_validator, code_analyzer, code_editor | 0.500 |
| 14 | issue_reproducer, code_locator, bug_analyzer | 0.417 |
| 15 | test_validator, patch_editor, repo_explorer | 0.500 |
| 16 | code_locator, issue_reproducer, bug_analyzer | 0.455 |
| 17 | test_validator, patch_editor, codebase_explorer | 0.583 |
| 18 | issue_reproducer, bug_localizer, bug_analyzer | 0.250 |
| 19 | code_patcher, fix_validator, issue_analyzer | 0.417 |
| 20 | code_locator, reproducer, code_editor | 0.417 |
A.2.4 Manually Designed Subagent Configurations
A.3 Additional Experimental Results
A.3.1 Effect of Design-Set Size
We evaluated whether the size of the design set affects the performance of the discovered sub-agents and found no meaningful impact. We sample 6 unique problems from SWE-Bench-Verified and sample one problem from each to make a small design set of 6 problems, and sample two problems from each to make a large design set of 24 problems. Across different design-set sizes, the resulting sub-agents achieve similar performance as shown in Table 6.
| Design Set Size | Resolution Rate |
| 6 | 21% |
| 12 | 20% |
| 24 | 19% |
A.4 Additional Experimental Setup Details
A.4.1 Implementation Details for Evolutionary Baseline
For comparison with BOAD, we implemented an evolutionary multi-agent design baseline where the orchestrator and three sub-agent prompts are bundled together and treated as a single evolutionary individual. We use three sub-agents paired with an orchestrator because BOAD chooses three sub-agents when optimizing the sub-agents in the experiments (Section 5.2). At each iteration of the evolution, the system generates three new subagents sequentially, given the previous round subagent configurations and their measured helpfulness and success rates (using the same prompts as BOAD, provided in A.1.2). Next, both the sub-agent warm-up and the orchestrator construction follow the same procedures used in BOAD. The generated orchestrator and subagents are then run on the same design set to get the helpfulness and success rate for the next iteration.
A.4.2 API Cost Comparison with BOAD
The evolutionary baseline also requires Claude calls—both to propose new orchestrator/sub-agent prompts and to perform LLM-as-judge scoring. Under the same setup, each evolution iteration costs $2.33, whereas each BOAD iteration costs $0.96, indicating that evolution is substantially more expensive to run. Evolution incurs higher cost because it mutate the bundle of orchestrator and sub-agents at each iteration.
Appendix B The Usage of Large Language Models (LLMs)
In our work, we used large language models (LLMs) for two purposes: (1) as a general writing assistant to check the grammar of the manuscript for readability, and (2) as the model component of our proposed system BOAD, where LLMs function as the evolution engine, meta-/sub-agents, and evaluation judges in experiments with mainstream coding agents. All research ideas, methodological contributions, and conclusions are solely those of the authors, who take full responsibility for the content of this work.

### v2

@@
 Abstract
-Large language models (LLMs) have shown strong reasoning and coding capabilities, yet they struggle to generalize to real-world software engineering (SWE) problems that are long-horizon and out of distribution.
+Large language models (LLMs) have shown strong reasoning and coding capabilities, yet they struggle to generalize to real-world software engineering (SWE) problems that are long-horizon and out-of-distribution.
 Existing systems often rely on a single agent to handle the entire workflow—interpreting issues, navigating large codebases, and implementing fixes—within one reasoning chain.
@@
 On SWE-bench-Verified, our method consistently outperforms single-agent systems and manually designed multi-agent systems.
-On SWE-bench-Live [30], which includes more recently collected issues and presents out-of-distribution challenges, our system achieves 3rd place on the leaderboard using a 36B model—outperforming larger-scale systems based on Claude and GPT-4.
+On SWE-bench-Live [30], which includes more recently collected issues and presents out-of-distribution challenges, our system ranked second on the leaderboard at the time of evaluation using a 36B model—outperforming larger-scale systems based on Claude and GPT-4.
 To our best knowledge, our work is the first to show generalization improvements using automatically discovered hierarchical multi-agent systems on challenging long-horizon interactive tasks like SWE-bench.
@@
 Temporal Abstraction via Semi-MDP (SMDP)
-Semi-Markov Decision Process (SMDP) framework [39] is widely used to mitigate long-horizon sparse reward problems.
+The Semi-Markov Decision Process (SMDP) framework [39] is widely used to mitigate long-horizon sparse reward problems.
 Instead of issuing primitive actions , the orchestrator selects a temporally extended action (option) .
-Each option corresponds to a sub-agent that executes a sequence of actions until termination, after which control returns at , where denotes the duration of an sub-agent.
+Each option corresponds to a sub-agent that executes a sequence of actions until termination, after which control returns at , where denotes the duration of a sub-agent.
 This reduces decision frequency and simplifies planning.
@@
 Implementation details for the evolutionary search are provided in Appendix A.4.1.
-When using the same number of evaluation instances (i.e number of SWE-Bench patches generated), we find that the sub-agents generated by the evolutionary search achieve worse performance (17.0% vs 20.0% on SWE-Bench-Live) than BOAD.
-Additionally, for the same number of iterations, the cost of Claude API calls is more than double than that of BOAD due to the need of generating many sub-agents at each iteration, whereas BOAD reuses subagents across iterations.
+When using the same number of evaluation instances (i.e., number of SWE-Bench patches generated), we find that the sub-agents generated by the evolutionary search achieve worse performance (17.0% vs 20.0% on SWE-Bench-Live) than BOAD.
+Additionally, for the same number of iterations, the cost of Claude API calls is more than double that of BOAD due to the need of generating many sub-agents at each iteration, whereas BOAD reuses subagents across iterations.
 | Scale | Model | Scaffold | Verified Resolved (%) | Live Resolved (%) |
@@
 When the fix required edits at multiple call sites or modules, the single agent often either over-edited unrelated regions or missed one or more necessary locations.
-The hierarchical system mitigated both omission and extraneous edits by delegating to a sub-agent for localization, which did a thorough analysis of the repository before making any targetted modifications.
+The hierarchical system mitigated both omission and extraneous edits by delegating to a sub-agent for localization, which did a thorough analysis of the repository before making any targeted modifications.
 3.
@@
 Inspecting these outputs, we found that erroneous sub-agent outputs (e.g., incomplete span identification, misinterpretation of the issue) were accepted as ground truth by the orchestrator, leading subsequent steps astray.
-Because there is no intermediate validation or self-checking, the orchestrator has limited ability to recover from such upstream.
+Because there is no intermediate validation or self-checking, the orchestrator has limited ability to recover from such upstream errors.
 These observations align with our hypothesis: hierarchical delegation constrains edit scope and improves coverage of multi-site fixes, but introduces a new dependency on the quality of sub-agent handoffs.
@@
 A.1 Prompt templates
-A.1.1 Prompt template for refining subagent during warmup stage
+A.1.1 Prompt template for refining sub-agent during warmup stage
 A.1.2 Meta agent prompts
 A.1.3 Custom orchestrator plan prompt
-A.1.4 Prompt template for checking if a subagent was helpful in a given instance
+A.1.4 Prompt template for checking if a sub-agent was helpful in a given instance
 A.1.5 Orchestrator-only Prompt
-A.2 Subagents
-A.2.1 BOAD Top 2 Discovered Subagent Configurations for Seed-OSS-36B-Instruct
-A.2.2 All BOAD Discovered Subagents
+A.2 Sub-agents
+A.2.1 BOAD Top 2 Discovered Sub-agent Configurations for Seed-OSS-36B-Instruct
+A.2.2 All BOAD Discovered Sub-agents
 Here we report all sub-agents discovered over 100 iterations of BOAD.
 For each sub-agent, we report the iteration number where it was first proposed (Generated Iteration), the number of times it was selected during optimization (), and its final hindsight helpfulness rate.
-| Subagent | Generated Iteration | Helpfulness | |
+| Sub-agent | Generated Iteration | Helpfulness | |
 | code_navigator | 1 | 1140 | 0.933 |
@@
 | refactor_architect | 96 | 24 | 0.042 |
-A.2.3 All Evolution Discovered Subagents
+A.2.3 All Evolution Discovered Sub-agents
 Here we report all bundles discovered over 20 iterations of BOAD (one for each iteration).
-For each sub-agent, we report its iteration number, the generated subagents, and the bundle’s success rate.
+For each sub-agent, we report its iteration number, the generated sub-agents, and the bundle’s success rate.
 Note that sub-agents with the same name across different iterations may have different system/instance templates because sub-agents are not reused.
@@
 This explains why code-editing sub-agents like code_fixer do not exhibit high average helpfulness.
-| Iteration | Subagents | Success Rate |
+| Iteration | Sub-agents | Success Rate |
 | 1 | issue_localizer, issue_reproducer, patch_editor | 0.333 |
@@
 | 20 | code_locator, reproducer, code_editor | 0.417 |
-A.2.4 Manually Designed Subagent Configurations
+A.2.4 Manually Designed Sub-agent Configurations
 A.3 Additional Experimental Results
@@
 We use three sub-agents paired with an orchestrator because BOAD chooses three sub-agents when optimizing the sub-agents in the experiments (Section 5.2).
-At each iteration of the evolution, the system generates three new subagents sequentially, given the previous round subagent configurations and their measured helpfulness and success rates (using the same prompts as BOAD, provided in A.1.2).
+At each iteration of the evolution, the system generates three new sub-agents sequentially, given the previous round sub-agent configurations and their measured helpfulness and success rates (using the same prompts as BOAD, provided in A.1.2).
 Next, both the sub-agent warm-up and the orchestrator construction follow the same procedures used in BOAD.
-The generated orchestrator and subagents are then run on the same design set to get the helpfulness and success rate for the next iteration.
+The generated orchestrator and sub-agents are then run on the same design set to get the helpfulness and success rate for the next iteration.
 A.4.2 API Cost Comparison with BOAD
@@
 Under the same setup, each evolution iteration costs $2.33, whereas each BOAD iteration costs $0.96, indicating that evolution is substantially more expensive to run.
-Evolution incurs higher cost because it mutate the bundle of orchestrator and sub-agents at each iteration.
+Evolution incurs higher cost because it mutates the bundle of orchestrator and sub-agents at each iteration.
 Appendix B The Usage of Large Language Models (LLMs)